Prove that an open interval in R is an open set and a closed intervals is a closed set
1
Expert's answer
2020-05-19T18:56:21-0400

An open set "S" in "R" is such set, that if "x \in S" then there exist such "\varepsilon >0" , that "(x-\varepsilon, x+\varepsilon)" "\in S". Since for every "x \in (a, b)" , "a < x<b", then (if either "a\neq -\infin" or "b \neq + \infin") we will define "\varepsilon" as "min(|x-a|, |x-b|)>0."

If, for instance, "a = - \infin" and "b \neq +\infin", then "\varepsilon := |x-b|>0", same for "a \neq - \infin, b = +\infin". Notice, that "a" can not be equal to "+\infin", as well as "b" can not be equal to "+ \infin", since "a<b". If "a= -\infin, b = + \infin" simultaneously, we let "\varepsilon = 1".

With such "\varepsilon", by the definition, "(x- \varepsilon, x+\varepsilon) \in (a, b)," if "x \in (a, b)". Thus, "(a, b)" is an open set.


A closed set "S" in "R" is such set, that its complement is open. Let's show that "[a, b]^{C}" is open. Indeed, "[a, b]^{C} = (-\infin,a) \cup (b, +\infin)". Notice, that If either "a = -\infin", or "b = + \infin", then there is only one interval in the complement.

"[a, b]" is the set of following elements: "x \in R", "a \le x \le b" . We have already proved that open interval is an open set. Thus, since the union of open sets is also an open set, we obtan that "(-\infty,a)\cup (b, +\infty)" is an open set too. (if there was only one interval in the complement, then the last sentence can be ommitted).

So, "[a, b]"  is a closed set.